in the the lab experiment of Friedel-Crafts Alkylation: 0.4 ml of benzene was mixed with 0.8 ml of 2-chloro 2 methyl propane (the ratio is 1 to 2) to produce 1-4 dibutyl benzene. how to calculate percent yeild and what is the limiting reagent?
Answer #1

Determination of limiting reagent :

Number of moles of benzene =

weight(grams)/molecular weight

Weight of benzene = volume* density

= 0.4 ml * 0.876

= 0.3504 grams

Moles of benzene = 0.3504/78

= 0.00449 moles (or) 4.49 m.moles.

Number of moles of 2-chloro-2-methyl propane = weight/molecular weight

Weight of 2-chloro-2-methylpropane = density * volume

= 0.851*0.8

= 0.6808 grams.

Moles of 2-chloro-2-methylpropane = 0.6808/92.57

= 0.007354 moles (or) 7.3544 m.moles.

Among the two reactants benzene has less number of moles.thus benzene is the limiting reagent.

Calculation of theoretical yield:

78 grams of benzene produces 190.32 grams of 1,4-di butyl benzene.

0.3504 grams of benzene produces ? grams of 1,4- diDbutyl benzene

= 0.3504*190.32/78

= 0.854 grams.

Theoretical yield (100% yield) = 0.854 ki.

% of yield = experimental weight (grams)*100/theoretical yield

But experimental weight is not given in the data.